OpenStudy (anonymous):

How do I solve ln(-x) +ln3=ln(2x-15)

OpenStudy (anonymous):

use this property to combine the ln's on the left side of the equation: \[\ln(a)+\ln(b) = \ln(a*b)\]

OpenStudy (anonymous):

Thus the left side of your equation becomes: \[\ln(-x)+\ln(3)=\ln(-3x)\] Now that your equation look like this: \[\ln(-3x)= \ln(2x-15)\] get rid of the ln's, and you have: -3x = 2x - 15 -5x = -15 x = 3

OpenStudy (anonymous):

Just not sure if there is an actual solution since solving for 3 ends up with ln of a negative number on both sides.

myininaya (myininaya):

there is no real solution